OpenStudy (anonymous):

Please help?????????? simplify 9√40

OpenStudy (anonymous):

\[\sqrt{40} = \sqrt{4 * 10}\] \[\sqrt{40} = 2\sqrt{10}\] 9 * 2 = 18 Answer = 18sqrt(10)

OpenStudy (anonymous):

youre great...wow i shouldve known that! so simple!